Supposed a truck, 10 - m long travels at a speed of 2.5 x 108 m/s relative to two observers, Ana and Marie. Both observers have a can of orange spray paint, which they fire simultaneously when the respective ends of the truck are directly in front of them. When the truck stops, Ana,


Marie and the truck driver look at the orange paint spots. How far apart are these spots? How would the truck driver interpret the same event while in the moving truck?

1
Expert's answer
2022-02-06T14:32:30-0500

"l=l_0\sqrt{1-\frac{v^2}{c^2}}=5.5~m."